### v1

@@ 3.2 Causal Graph Construction and Updating
If new data suggests revised causal evidence for any edge, a causal mining algorithm (e.g., PC [15, 13]) is adopted to update edge weights. Let $S_{\text{new}}(u,v)$ be the new strength for $(u,r,v)$. If $S_{\text{new}}(u,v)\geq\theta$, that edge is added or revised in $\mathcal{G}_{C}$:
@@
3.3 Chain-of-Thought–Driven Causal Path Retrieval
@@
For each query, the LLM is prompted to produce a chain-of-thought (CoT) in short segments, each separated by $\rightarrow$” and ending with a numeric confidence in $[0,100]$. The chain is then split on these separators (omitting the final confidence) into segments $\{s_{1},s_{2},\dots,s_{T}\}$, each $s_{i}$ being a concise statement or state. Figure 2 illustrates this step-by-step retrieval process, showing how each segment is used to form intermediate sub-queries.
Figure 2: An example of chain-of-thought–driven retrieval. Each CoT segment triggers a partial query to the causal subgraph, producing candidate paths that are later fused.
@@ 3.4.2 Second-Stage Enhancement: Re-Injecting Paths and CoT
After selecting the best paths, they—together with the original CoT—are re-injected into the LLM for a final consistency check. During this phase, the model can cross-verify domain knowledge, remove spurious or contradictory branches, and produce a succinct “enhanced” summary to guide the final answer. This second-stage approach typically clarifies ambiguous steps or fuses partial sub-graphs into a coherent explanation. The final answer thus leverages both the curated paths and the chain-of-thought, improving coherence and interpretability.
Implementation Notes.
@@ 4 Experiments
This section evaluates the causal-first Graph RAG framework on two medical multiple-choice QA datasets derived from MedMCQA and MedQA. Both datasets contain clinically oriented questions across multiple medical subtopics. Only items mapping into the causal subgraph are retained for testing; each system is compared using Precision, Recall, and F1 Score as the evaluation metrics.
@@ 4.1 Experimental Setup
All retrievals occur against a SemMedDB-based knowledge graph (KG)***https://lhncbc.nlm.nih.gov/ii/semmeddb/, containing medically oriented entities (e.g., diseases, drugs, pathways) and relations (e.g., CAUSES, MANIFESTATION_OF, ASSOCIATED_WITH). The causal subgraph is constructed as described in Section 3.2, discarding edges whose cause-effect weight is below threshold $\theta$. Some queries that fail to match any node in this subgraph are excluded.
@@ 4.2 Experimental Results
As shown in Table 1, the LLMs leveraging the proposed method, which includes cause-effect filtering, CoT-aligned retrieval, and multi-stage path refinement, outperform direct single-step model responses across all three metrics—Precision, Recall, and Recall F1 Score. The results highlight that the proposed approach, particularly when applied to CGMT(GPT-4o mini), provides significant improvements over direct GPT-4o mini (with precision up to 10% absolute), showcasing the effectiveness of this method in enhancing model performance for knowledge-intensive tasks.
@@
In the second experiment, this research compare the proposed CGMT method with traditional Graph-RAG models. As shown in Table 1, the CGMT models, particularly CGMT(GPT-4o), consistently outperform the traditional Graph-RAG models across both datasets. For example, CGMT(GPT-4o) achieves a precision of 92.90 on MedMCQA and XXX on MedQA, significantly higher than the corresponding values for Graph-RAG(GPT-4o) (86.88 and XXX). This demonstrates that the proposed CGMT method, with its integrated cause-effect filtering and CoT-aligned retrieval, offers substantial improvements over traditional Graph-RAG models, especially in handling complex knowledge-intensive tasks.
Table 1: Evaluation results on the MedMCQA dataset: comparing CGMT (Enhanced) with Graph-RAG (Normal_rag) and direct responses (Derelict) under different model variants.
Model Accuracy(%) Recall(%) F1 Score
CGMT (GPT-4o) 92.90 93.33 0.95
CGMT (GPT-4) 87.98 88.57 0.92
CGMT (GPT-4o mini) 82.51 82.67 0.89
Graph-RAG (GPT-4o) 86.89 86.81 0.91
Graph-RAG (GPT-4) 82.51 83.56 0.88
Graph-RAG (GPT-4o mini) 75.41 78.91 0.84
GPT-4o 86.89 86.30 0.91
GPT-4 84.15 84.83 0.89
GPT-4o mini 72.13 76.51 0.82
Table 2: Ablation Study for CGMT Method with Various Enhancements on the MedMCQA dataset.
Method GPT-4o(%) GPT-4(%) GPT-4o mini(%)
CGMT (Enhanced) 92.90 87.98 82.51
KG-only 91.80 85.25 79.78
Remove LLM Enhanced 90.16 86.89 79.78
Remove Enhancer 90.71 85.25 79.78
@@ 4.3 Chain-of-Thought Cross-Model Evaluation
Cross-model scenarios were tested, where one model’s Chain-of-Thought (CoT) is used by another. These tests involve three stages: the first stage is CoT generation, where one model generates the CoT; the second stage is enhancement, where the CoT is refined and enhanced by a different model; and the third stage is inference, where the final answer is generated based on the enhanced CoT. Table 3 details the results of three such trials across these stages. Baseline references ( 92.90% for GPT-4o, 83.06% for GPT-4o mini ) come from the “CGMT” row in Table 1.
@@
Overall, the findings confirm that causal prioritization and CoT alignment substantially enhance retrieval-augmented medical QA. Targeting cause-effect edges narrows the KG to more consistent subpaths, while segmenting retrieval according to the chain-of-thought preserves logical clarity. Each LLM—GPT-4o, GPT-4, or GPT-4o-mini—yields its best accuracy under “enhanced,” underscoring the broad applicability of this approach to different model scales. Moreover, cross-model CoT trials underscore that reasoning trace quality significantly influences final performance, particularly for complex, knowledge-intensive queries.
@@ 5 Discussion
Although the experiments indicate that a causal-first Graph RAG framework, aligned with chain-of-thought (CoT) prompting, can notably enhance multi-hop medical QA, certain limitations and open questions remain:
@@
Despite leveraging a large SemMedDB-based subgraph, certain clinically relevant edges may be missing, forcing fallback retrieval from correlation-based links. The wide connectivity and shallow semantics (e.g., “related to”) can also cause clutter. More robust causal-mining and curated domain expansions could mitigate these drawbacks.
@@
The results validate the advantages of aligning causal-first retrieval with CoT. Further gains may be realized by stabilizing CoT prompts (via self-consistency sampling or advanced prompt engineering) and refining the knowledge graph itself (via domain-specific expansions). As LLMs and causal graphs continue to improve, uniting explicit cause-effect edges with chain-of-thought reasoning offers a compelling paradigm for knowledge-intensive domains.
@@ 6 Conclusion
This work has presented a comprehensive framework integrating causal-first knowledge graph filtration with chain-of-thought (CoT) guided retrieval, culminating in a multi-stage path enhancement strategy. Departing from correlation-based Graph RAG, the method prioritizes cause-effect edges in a large, densely connected medical KG and aligns retrieval steps with the model’s CoT prompts by segment. Empirical evaluation on medical QA tasks (MedMCQA and MedQA) shows consistent accuracy gains across three LLMs (GPT-4o, GPT-4, GPT-4o-mini). Ablations verify that each component, from causal graph construction to second-stage LLM refinement, contributes to final performance. Cross-model CoT experiments highlight that reasoning trace quality strongly determines downstream success.The key takeaways from this work are as follows:
@@
  • Chain-of-Thought Alignment: Dividing queries into short, single-state steps (using $\rightarrow$”) and matching each with a subgraph query fosters accurate multi-hop solutions.
@@
Potential extensions include refining causal-mining algorithms to improve edge quality, using more robust entity-linking to reduce retrieval misses, and incorporating advanced domain modules (e.g., biomedical reading comprehension). Furthermore, extending these methods to legal or scientific literature QA might similarly profit from explicit causal priorities and CoT-driven retrieval. As knowledge graphs and LLMs continue to evolve, this synergy between cause-effect edges and structured chain-of-thought reasoning is likely to become increasingly pivotal.

### v2

@@ 3.2 Causal Graph Construction and Updating
If new data suggests revised causal evidence for any edge, a causal mining algorithm (e.g., PC [14]) is adopted to update edge weights. Let $S_{\text{new}}(u,v)$ be the new strength for $(u,r,v)$. If $S_{\text{new}}(u,v)\geq\theta$, that edge is added or revised in $\mathcal{G}_{C}$:
@@
3.3 Chain-of-Thought-Driven Causal Path Retrieval
This section outlines a method for causal path retrieval driven by the chain-of-thought process, which facilitates more coherent and targeted reasoning, as shown in Figure 2.
@@ 3.3 Chain-of-Thought-Driven Causal Path Retrieval
For each query, the LLM is prompted to produce a chain-of-thought (CoT) in short segments, each separated by $\rightarrow$" and ending with a numeric confidence in $[0,100]$. The chain is then split on these separators (omitting the final confidence) into segments $\{s_{1},s_{2},\dots,s_{T}\}$, each $s_{i}$ being a concise statement or state.
Figure 2: An example of chain-of-thought-driven retrieval. Each CoT segment triggers a partial query to the causal subgraph, producing candidate paths that are later fused.
@@ 3.4.2 Second-Stage Enhancement: Re-Injecting Paths and CoT
After selecting the best paths, they, together with the original CoT, are re-injected into the LLM for a final consistency check. During this phase, the model can cross-verify domain knowledge, remove spurious or contradictory branches, and produce a succinct "enhanced" summary to guide the final answer. This second-stage approach typically clarifies ambiguous steps or fuses partial sub-graphs into a coherent explanation. The final answer thus leverages both the curated paths and the chain-of-thought, improving coherence and interpretability.
Implementation Notes
@@ 4 Experiments
This section evaluates the causal-first Graph-RAG framework on two medical multiple-choice QA datasets derived from MedMCQA[9] and MedQA[6]. Both datasets contain clinically oriented questions across multiple medical subtopics. Only items mapping into the causal subgraph are retained for testing; each system is compared using Precision, Recall, and F1 Score as the evaluation metrics.
@@ 4.1 Experimental Setup
All retrievals occur against a SemMedDB-based knowledge graph (KG)***https://lhncbc.nlm.nih.gov/ii/semmeddb/, containing medically oriented entities. The causal subgraph is constructed as described in Section 3.2, discarding edges whose cause-effect weight is below threshold $\theta$. Some queries that fail to match any node in this subgraph are excluded.
@@ 4.2 Experimental Results
As shown in Table 1, the LLMs leveraging the proposed method, which includes cause-effect filtering, CoT-aligned retrieval, and multi-stage path refinement, outperform direct single-step model responses across all three—Precision, Recall, and F1 Score. The results highlight that the proposed approach, particularly when applied to CGMT(GPT-4o mini), provides significant improvements over direct GPT-4o mini (with precision up to 10% absolute), showcasing the effectiveness of this method in enhancing model performance for knowledge-intensive tasks.
@@
In the second experiment, this research compare the proposed CGMT method with traditional Graph-RAG models. As shown in Table 1, the CGMT models, particularly CGMT(GPT-4o), consistently outperform the traditional Graph-RAG models across both datasets. For example, CGMT(GPT-4o) achieves a precision of 92.90 on MedMCQA and 88.36 on MedQA, significantly higher than the corresponding values for Graph-RAG(GPT-4o) (86.89 and 83.07). This demonstrates that the proposed CGMT method, with its integrated cause-effect filtering and CoT-aligned retrieval, offers substantial improvements over traditional Graph-RAG models, especially in handling complex knowledge-intensive tasks.
@@ 4.2.3 Experiment 3: Ablation Study
Table 1: Evaluation results for LLMs using the proposed method compared with traditional Graph-rag and direct responses across two datasets.
Model MedMCQA MedQA
Precision Recall F1 Score Precision Recall F1 Score
CGMT(GPT-4o) 92.90 93.33 0.95 88.36 74.03 0.84
CGMT(GPT-4) 87.98 88.57 0.92 74.07 55.21 0.68
CGMT(GPT-4o mini) 82.51 82.67 0.89 72.49 53.68 0.66
Graph-RAG(GPT-4o) 86.89 86.81 0.91 83.07 70.37 0.81
Graph-RAG(GPT-4) 82.51 83.56 0.88 72.49 53.47 0.68
Graph-RAG(GPT-4o mini) 75.41 78.91 0.84 62.96 44.86 0.58
GPT-4o 85.52 86.30 0.90 81.48 65.52 0.78
GPT-4 84.15 84.83 0.89 69.84 51.00 0.64
GPT-4o mini 72.13 76.51 0.82 60.85 43.36 0.57
Table 2: Ablation Study for CGMT Method with Various Enhancements Across Two Datasets.
Model Precision of MedMCQA(%) Precision of MedQA(%)
GPT-4o GPT-4 GPT-4o mini GPT-4o GPT-4 GPT-4o mini
CGMT 92.90 87.98 83.06 88.36 74.07 72.49
KG-only (Knowledge Graph) 91.80 85.25 79.78 87.30 70.37 70.90
Remove LLM Enhanced 90.16 86.89 81.42 85.18 68.76 68.79
Remove Enhancer 90.71 83.61 78.14 86.24 71.96 69.84
@@ 4.3 Chain-of-Thought Cross-Model Evaluation
Cross-model scenarios were tested, where one model’s Chain-of-Thought (CoT) is used by another. These tests involve three stages: the first stage is CoT generation, where one model generates the CoT; the second stage is enhancement, where the CoT is refined and enhanced by a different model; and the third stage is inference, where the final answer is generated based on the enhanced CoT. Table 3 details the results of three such trials across these stages. Baseline references come from the “CGMT” row in Table 1.
@@ 5 Discussion
Although the experiments indicate that a causal-first Graph-RAG framework, aligned with chain-of-thought (CoT) prompting, can notably enhance knowledge-intensive reasoning tasks, certain limitations and open questions remain:
@@
Despite leveraging a large SemMedDB-based subgraph, certain clinically relevant edges may be missing, forcing fallback retrieval from correlation-based links. The wide connectivity and shallow semantics (e.g., "related to") can also cause clutter. More robust causal-mining and curated domain expansions could mitigate these drawbacks.
@@
The results validate the advantages of aligning causal-first retrieval with CoT. Further gains may be realized by stabilizing CoT prompts (via self-consistency sampling or advanced prompt engineering) and refining the knowledge graph itself (via domain-specific expansions).
@@ 6 Conclusion
This work has presented a comprehensive framework integrating causal-first knowledge graph filtration with chain-of-thought (CoT) guided retrieval, culminating in a multi-stage path enhancement strategy. Departing from correlation-based Graph-RAG, the method prioritizes cause-effect edges in a large, densely connected medical KG and aligns retrieval steps with the model’s CoT prompts by segment. Empirical evaluation on medical QA tasks (MedMCQA and MedQA) shows consistent accuracy gains across three LLMs (GPT-4o, GPT-4, GPT-4o-mini). Ablations verify that each component, from causal graph construction to second-stage LLM refinement, contributes to final performance. Cross-model CoT experiments highlight that reasoning trace quality strongly determines downstream success. The key takeaways from this work are as follows:
@@
  • Chain-of-Thought Alignment: Dividing queries into short, single-state steps (using $\rightarrow$") and matching each with a subgraph query fosters accurate multi-hop solutions.
@@
Potential extensions include refining causal-mining algorithms, improving entity-linking to reduce retrieval misses, and integrating advanced domain modules. Additionally, applying these methods to legal or scientific literature QA could benefit from causal prioritization and CoT-driven retrieval. As knowledge graphs and LLMs evolve, the synergy between cause-effect edges and structured reasoning will become increasingly important.